OpenStudy (anonymous):

Help Pleasee.!! Marjorie drives 17.3 miles each way to work at an average of $0.20 per mile. She pays $3.25 each day to park and works an average of 22 days per month. If she joins a car pool with two other workers, what will her average monthly cost be?

OpenStudy (anonymous):

17.3 + 17.3 = 34.6 (this is how much she drives daily. 34.6 x 22 = 761.20 (miles driven per month) 761.20 x .20 = 152.24 (money used for gas for the month) 22 x 3.25 = 71.50 (how much she pays for parking for the month. for a total cost of (152.24 + 71.50) = 223.74 If there is 3 people, including Marjorie, it would have to be divided by 3 because each person is supposed to pay for 1/3 of the total cost. 223.74/3 = 74.58 Each person will pay 74.58 Do you understand what I just did ?

OpenStudy (anonymous):

yeaa :) thanks

OpenStudy (anonymous):

your quite welcome...that is what I am here for :)

OpenStudy (anonymous):

how did you get the 22for the miles driven per month?

OpenStudy (anonymous):

22 is the number of days she works and she drives 34.6 miles per day. So I multiplied 22 by 34.6 to determine the number of miles driven per month.

OpenStudy (anonymous):

oo, can you help me w another one?

OpenStudy (anonymous):

possibly...what you got ?

OpenStudy (anonymous):

Les drives 4 miles each way to work at an average of $0.195 per mile. He pays $2.25 a day to park and works an average of 24 days per month. What is his average monthly cost?

OpenStudy (anonymous):

sorry itsz cux oo dont understand consumers math

OpenStudy (anonymous):

4 miles each way to work means he drives 8 miles per day. He works 24 days so he drives (8 x 24) = 192 (this is how many miles per month he drives. It costs him (in gas) 192 x 0.195 = 37.44 (cost of gas per month) He pays 2.25 a day for parking and he works 24 days, so he pays 24 x 2.25 = 54.00 per month for parking. Now we add... gas : 37.44 parking : 54.00 -------------add total = 91.44 <==average monthly cost

OpenStudy (anonymous):

okai so my next questions is the same numbers but it says ; Les drives 4 miles each way to work at an average of $0.195 per mile. He pays $2.25 a day to park and works an average of 24 days per month. If Les joins a car pool with three other workers, what will his average monthly cost be?

OpenStudy (anonymous):

If average monthly cost for 1 person is 91.44, then the average monthly cost for 3 people would be 91.44/3 = 30.48

OpenStudy (anonymous):

but by answers are only, $22.86, $ 21.56, & $18.18

OpenStudy (anonymous):

did I get the first part of the question correct....the one where I answered 91.44

OpenStudy (anonymous):

yes thats on my list

OpenStudy (anonymous):

oh wait....he joins with 3 other people.....that makes a total of 4 people....so it would be 91.44/4 = 22.86

OpenStudy (anonymous):

ooo, ii See :)

OpenStudy (anonymous):

the next questions says she dirves 5.5 miles each way at a cost of &0.26

OpenStudy (anonymous):

per mile and has no parking fee. On an average month of 21 days, what is her monthly cost?

OpenStudy (anonymous):

do i multiply 5.5 twice?

OpenStudy (anonymous):

yes...because it is 5.5 miles there and 5.5 miles back.....that will tell you how many miles he drive per day

OpenStudy (anonymous):

then multiply it by .26?

OpenStudy (anonymous):

ooo & Jeri is in a car pool with one other worker

OpenStudy (anonymous):

...well you can to determine how much she pays for gas each day, or you can multiply it by 21 to determine how many miles per month then multiply that by .26. It can be done either way. Remember if you multiply the miles per day by .26, that is only telling you the cost per day, you would still have to multiply by 21 for the cost per month

OpenStudy (anonymous):

Even if he is in a carpool with another person...don't worry about that now. Figure out what 1 person pays then divide that by 2

OpenStudy (anonymous):

wait . im lost ?

OpenStudy (anonymous):

ok.....there is two ways to do this.... First way : 5.5 x 2 = 11 (miles per day) 21 x 11 = 231 (miles per month) 231 x .26 = 66.06 <==monthly cost or this way... 5.5 x 2 = 11 11 x .26 = 2.86 (cost of gas per day) 21 x 2.86 = 60.06 (cost of gas per month) Either way it cost 1 person 60.06 per month. For two people it would cost 60.06/2 = 30.03 per person

OpenStudy (anonymous):

ooooo, now ii get it

OpenStudy (anonymous):

what does a car pool mean cuz my

OpenStudy (anonymous):

next question has that word

OpenStudy (anonymous):

a carpool means there is more then one person traveling in the car and you split the cost with them.

OpenStudy (anonymous):

Jeri is in a car pool with one other worker. She drives 5.5 miles each way at a cost of $0.26 per mile and has no parking fee. On an average month she works 21 days. If Jeri decides to drive alone, how much more will it cost her per month?

OpenStudy (anonymous):

is it like the one u just did?

OpenStudy (anonymous):

yes....figure it out for one person , then figure it out for 2 people

OpenStudy (anonymous):

so its the same answer isnt it?

OpenStudy (anonymous):

not exactly....for 1 person it cost 60.06, for 2 people it cost 30.03. It is asking how much more will it cost subtract 60.06 - 30.03 = 30.03 It would cost Jeri 30.03 more per month to drive alone then it would for a carpool.

OpenStudy (anonymous):

It is the same answer...lol

OpenStudy (anonymous):

oooo, thats why ii thought it was the ssame

OpenStudy (anonymous):

Are you understanding any of this ?

OpenStudy (anonymous):

yea i am thanks alot :)

OpenStudy (anonymous):

your welcome...I am glad to help :)

OpenStudy (anonymous):

for this question i would use car pool riqht? Workers form a ________ to share the costs of driving to work.

OpenStudy (anonymous):

yes... carpool is when you share a ride with 1 or more people where you divide up the costs equally.

OpenStudy (anonymous):

oo, for this one is it time zone? To plan your travel itinerary, you use________.

OpenStudy (anonymous):

probably not time zone...Itinerary is a detailed plan for a journey. Is a map one of the answers ?

OpenStudy (anonymous):

so schedules?

OpenStudy (anonymous):

that would work :)

OpenStudy (anonymous):

oo okaii

OpenStudy (anonymous):

do u kno how to dived with fractions?

OpenStudy (anonymous):

yes...I do :)

OpenStudy (anonymous):

give me a problem and I will show you

OpenStudy (anonymous):

how do u do, 4 1/2 x 5

OpenStudy (anonymous):

its not even a fraction or is it?

OpenStudy (anonymous):

thats not division...did you type it correctly ? I can show you how to do this problem...is it typed correctly ?

OpenStudy (anonymous):

4 1/2 × 5 yea dats how its on my page

OpenStudy (anonymous):

do you know how to turn mixed numbers into improper fractions ?

OpenStudy (anonymous):

if its multipication u divide dont u?

OpenStudy (anonymous):

no...if its multiplication you multiply, however, if its division you end up multiplying as well.

OpenStudy (anonymous):

To do this problem you have to turn the mixed number into an improper fraction. Do you know how to do that ?

OpenStudy (anonymous):

nope :(

OpenStudy (anonymous):

let me teach you.... 4 1/2 is a mixed number because it has a whole number and a fraction. To turn 4 1/2 into an improper fraction you take the whole number (4)and multiply it by the denominator (2) which equals 8. Then take the 8 and add it to the numerator (1) which equals 9. Then you put the 9 over the original denominator (2) so it would be 9/2. 4 1/2 turned to an improper fraction is 9/2. Do you understand this ?

OpenStudy (anonymous):

oo okai now ii member this,'

OpenStudy (anonymous):

continuing with the problem.. 4 1/2 x 5 = 9/2 x 5/1 = 45/2 = 22 1/2 Do you understand ?

OpenStudy (anonymous):

omg nopee

OpenStudy (anonymous):

nvm ii do get it now

OpenStudy (anonymous):

aa nvm i dont get it i tryed w the next question and i didnt get a whole num ber

OpenStudy (anonymous):

You can turn improper fractions into whole numbers. That is how I got a whole number. I originally got 45/2 which when turned to a mixed number is 22 1/2

OpenStudy (anonymous):

and who do u do do it w money? 12 x $3.01

OpenStudy (anonymous):

just multiply...12 x 3.01 = 36.12.....no fractions involved.

OpenStudy (anonymous):

oo okii

OpenStudy (anonymous):

do you have anymore questions ?

OpenStudy (anonymous):

ya but this one is way difrent way diffrent,

OpenStudy (anonymous):

how different ?

OpenStudy (anonymous):

Name: Gregory Allen Destination: New Orleans, LA Fare type: coach Cost per mile:________

OpenStudy (anonymous):

thats the question

OpenStudy (anonymous):

let me look at it and see what I can do

OpenStudy (anonymous):

okai thanks

OpenStudy (anonymous):

It will take Gregory Allen 1434 miles.It will cost him $168 for 1434 miles. So you want to know how much per mile it would cost. You divide 1434 by 168 and that will give you the cost per mile.

OpenStudy (anonymous):

wow. i got 8.53571.....

OpenStudy (anonymous):

my answers are $0.02, 0.12, $ 0.22

OpenStudy (anonymous):

let me try this again.... 1434/168 = 1/x (1434 miles to $168 = 1 mile to x $) now cross multiply... (1434)(x) = (168)(1) 1434x = 168 x = 168/1434 = .117 rounded to .12 sorry about that...I messed up on the first explanation.

OpenStudy (anonymous):

wait how u get x?

OpenStudy (anonymous):

x stands for cost per mile. This is what we are looking for. This is our unknown variable

OpenStudy (anonymous):

oo

OpenStudy (anonymous):

x is what I just solved for.

OpenStudy (anonymous):

why do i keep gettinq decimal w long numbers?

OpenStudy (anonymous):

it happens....sometimes it is better to round the number to make it shorter.

OpenStudy (anonymous):

look my

OpenStudy (anonymous):

huh ?

OpenStudy (anonymous):

next question is Name: Lola Wayne Destination: Seattle, WA Fare type: coach Cost of round-trip ticket:________

OpenStudy (anonymous):

do i divide 82 to 701

OpenStudy (anonymous):

round trip means there and back. Seattle is 701 miles one way, and costs $82 (one way) round trip is 1402 miles, and costs (82 x 2) = $164 total cost for round trip to Seattle is $164

OpenStudy (anonymous):

you do not divide, you multiply

OpenStudy (anonymous):

you are not looking for the cost per mile, you are looking for the total cost

OpenStudy (anonymous):

why is this so confsinqq

OpenStudy (anonymous):

look this one is diffrent

OpenStudy (anonymous):

Dennis is going from Atlanta to Los Angeles with a 45-minute stop in Salt Lake City. If his departure time from Atlanta is 8:45 A.M., what is his expected time of arrival in Los Angeles?